OpenStudy (maddytarin_98):

hat is the solution set of the quadratic inequality 6x^2+1 (less than or equal to sign) 0

OpenStudy (maddytarin_98):

*what

OpenStudy (maddytarin_98):

im missing information

OpenStudy (mathmate):

A sketch helps with all geometry questions and sometimes algebra questions!

OpenStudy (rolldextro):

There is no soluion.

OpenStudy (mathmate):

 From the graph, you see that the entire curve is above the x-axis, so there is no point on the curve that can be \(\le 0\).

OpenStudy (maddytarin_98):

yeah i get it now! thank you (: @mathmate

OpenStudy (mathmate):

What you need to show is that there is no solution in the real domain. \(6x^2+1 \le 0\) => \(x^2 \le -1/6\) Since \(x^2\ge\)0 for all x, so there is no solution for x\(\in\) R.